# Improving the Plausibility of Pressure Distributions Synthesized from Depth through Generative Modeling

### v1

@@ Abstract
Monitoring contact pressure in hospital beds is essential for preventing pressure ulcers and enabling real-time patient assessment. Current methods can predict pressure maps but often lack physical plausibility, limiting clinical reliability. This work proposes a framework that enhances plausibility via Informed Latent Space (ILS) and Weight Optimization Loss (WOL) with generative modeling to produce high-fidelity, physically consistent pressure estimates. This study also applies diffusion based conditional Brownian Bridge Diffusion Model (BBDM) and proposes training strategy for its latent counterpart Latent Brownian Bridge Diffusion Model (LBBDM) tailored for pressure synthesis in lying postures. Experiment results shows proposed method improves physical plausibility and performance over baselines: BBDM with ILS delivers highly detailed maps at higher computational cost and large inference time, whereas LBBDM provides faster inference with competitive performance. Overall, the approach supports non-invasive, vision-based, real-time patient monitoring in clinical environments.
@@ 1 Introduction
A major limitation of prior methods is their weak modeling of how pressure depends on human attributes (mass, height, gender). The proposed ILS addresses this by explicitly coupling these factors, producing more plausible and adaptive pressure maps. A practical advantage is flexibility: for the same subject’s depth image, adjusting mass, height, or gender yields distinct, physically consistent pressure distributions. This adaptability is clinically useful when patient characteristics change over time, since only the conditioning inputs need updating rather than retraining the model.
This study also introduces pressure synthesis via a conditional Brownian Bridge Diffusion Model (BBDM) and Latent Brownian Bridge Diffusion Model (LBBDM) [23]. We employed a two-stage training strategy for a conditional LBBDM, tailored for pressure synthesis. Additionally, a VQGAN model [42] trained on a distinct objective is integrated with the LBBDM for pressure generation. To see the impact of ILS in the denoising process, we have incorporated ILS into the denoising model.
We evaluate the proposed method on the publicly available Systematic Lying Postures (SLP) dataset [27] and compare it against established baselines, including BPBnet, BPWnet, and the AttnFnet + cGAN methods [8, 28].
@@ 2 Related Work
Conventionally, nurses reposition bedridden patients every two hours to reduce skin hydration, minimize prolonged mechanical load, and redistribute pressure [29]. As this repositioning frequency dates back to World War II [11], it remains a standard method for preventing pressure injuries caused by mechanical boundary conditions. However, studies by Peterson et al. [33] and Peterson et al. [32] suggest that the current repositioning protocol requires improvement, as poor pressure redistribution can still occur despite the scheduled turning.
@@
Building on these prior concepts, the proposed study introduces a novel approach for inferring plausible pressure distributions through conditional generative modeling. This study proposes the inclusion of anthropometric information to build interdependencies via ILS, and further improving pressure plausibility through WOL when predicting distributions in pixel space under a cGAN framework. Furthermore, in our work we synthesize pressure distributions using conditional BBDM, introduced pressure synthesis using BBDM + ILS, and introduces a two-stage training approach for a conditional LBBDM tailored for pressure synthesis. To the authors’ knowledge, no prior work has applied diffusion-based modeling for predicting pressure distributions.
@@
Although Betanet predict height and mass, this component remains independent of BPBnet and BPWnet, and does not establish any interrelationships between pressure and anthropometric attributes. Similarly, the approach proposed by Manavar et al. [28] lacks a mechanism to model such dependencies.
This study addresses these challenges by proposing an informed mechanism to incorporate plausibility into pressure inference through ILS and WOL. Additionally, ILS is integrated into the denoising process of the BBDM during training and into the pretrained network during LBBDM training to investigate its influence on pressure plausibility in generative denoising.
@@ 3.3 Training Strategy
This study employs conditional generative modeling approaches such as cGAN, BBDM, and LBBDM to synthesize pressure distributions and compare their performance, quality, and computational complexity trade-offs.
We have used encoder-decoder model architecture introduced by Manavar et al. [28] when training at the pixel level. For diffusion model training, we use OpenAI’s improved U-Net as the denoising model, consistent with the setup by Li et al. [23].
@@
Finally, we employed BBDM in the latent space to evaluate computational performance improvements alongside the quality of generated pressure distributions. The AttnFnet model was trained in a self-reconstruction manner using ILS to encode the relationships among mass, height, and gender during image reconstruction. Skip connections in the AttnFnet architecture were removed to minimize the influence of the encoder during the decoding process. To reduce the size of the denoising model, a convolutional layer was introduced after the bottleneck layer to decrease the feature channel dimensionality, followed by a second convolutional layer to restore the latent representation to its original feature channel dimensions (more in supplimentary Sec. 12).
@@
This study also utilized an unconditionally trained VQGAN [42] model on the CelebHQ dataset [21], in a manner consistent with the BBDM paper [23].
@@ 4 Evaluation
LBBDM proposed pretraining 0.6885 0.8687 1.0123 0.000309 35.5299
@@
Manavar et al. [28] AttnFnet 0.6580 0.8860 0.8014 0.000262 36.4592
Ours AttnFnet 0.6451 0.8795 0.6979 0.000270 36.1675
@@
LBBDM proposed pretraining 0.7017 0.8536 0.7269 0.000319 35.3887
Table 1: Test‑set comparison of MPPA, MSSIM, MFID, MSE, and MPSNR across methods at two training sample sizes.
All experiments were conducted using the dataset proposed by Liu et al. [27], and synthetic dataset by Clever [6]. The proposed methods were compared with baseline models from Clever et al. [8] and Manavar et al. [28].
@@ 4.1 Dataset
The Systematic Lying Postures (SLP) dataset [27] contains 102 healthy participants under three cover conditions and 45 unique poses, grouped into three sleeping postures (supine, left lateral, right lateral). Cover conditions include uncovered poses, poses covered by a 1 mm blanket, and poses covered by a 3 mm blanket.
@@ 4.2 Data Analysis
Proposed $\gamma=0.0182$ 0.6406 0.8814 0.8819 0.8013
@@
All models were evaluated under three different training conditions: (1) a sample size of $2745$ ($3\,\mathrm{k}$) using only uncovered postures from real data; (2) a sample size of $8235$ ($8\,\mathrm{k}$) including cover conditions from real data; and (3) a sample size of $203225$ ($203\,\mathrm{k}$) combining both real and synthetic data with all postures and covered conditions.
Standard evaluation metrics for 2D images were used to assess pressure generation quality: Mean Pixel Prediction Accuracy (MPPA) [28], Mean Structural Similarity Index (MSSIM) [40], Mean Fréchet Inception Distance (MFID) [37], MSE, and Mean Peak-Peak Signal-to-Noise Ratio (MPSNR) [13]. Fréchet inception distance (FID) was calculated with the original Inception-V3 model weights [16]. Table 1 presents metric scores for the proposed methods and baselines on the test dataset. In Tab. 1, BBDM with ILS outperformed all methods when trained on $8\,\mathrm{k}$ samples, while BBDM outperformed in $3\,\mathrm{k}$ uncovered samples.
When images were rescaled from the normalized space, MSE was recalculated in pressure space as shown in Tab. 2. Consistent trends were observed: the proposed AttnFnet method displayed increased MSE in pressure space when trained with $203\,\mathrm{k}$ samples, but error rate declined when setting $\gamma=0.0182$. Note that pressure ranges differ between real and synthetic data, and setting $\gamma=1.0$ prevents the model from generating valid samples.
@@
Participant BMs were calculated from actual and predicted pressure profiles using Eq. 9 and equations from [28]. Table 3 reports mean absolute body-mass differences among measured BMs ($M_{\text{meas}}$), BMs calculated from the ground truth pressures ($M_{\text{GT}}$), and from predicted pressures ($M_{\text{pred}}$). Further, BM plots in Fig. 4 enables detailed comparisons.
@@ 5 Ablation Study
Table 5 demonstrates that combining ILS and WOL yields more plausible pressure distributions with a reduction in MSE. Table 6 presents the effect of the WOL weight $\gamma$ on MSE and BM Mean Absolute Error (MAE): reducing $\gamma$ initially reduces the error but, degrades quality of calculated BM from predicted pressure distributions.
@@
Table 5: Effect of the ILS and WOL components on prediction error: MSE of pressure maps and BM MAE.
Method MSE ($kPa^{2}$) BM MAE
@@
Table 6: Effect of the WOL weight $\gamma$ on pressure MSE and BM MAE
Method MSE ($kPa^{2}$) BM MAE FID
@@
Table 7: Effect of diffusion sampling steps on pressure MSE, BM MAE, and FID;
@@ 6 Discussion and Future Work
The BBDM variants were not trained on the full $203\,\mathrm{k}$ samples due to memory and runtime constraints. By contrast, our method offers a stronger accuracy-plausibility trade-off with far lower memory use and shorter training time (see Tab. 3, Tab. 2, and supplimentary Sec. 9).
@@
With $8\,\mathrm{k}$ samples, BBDM+ILS does not reduce BM error (Tab. 3); but it improves the metrics, we observed when training on $3\,\mathrm{k}$ samples plausibility improves with ILS.
@@ Conclusion
This work predicts physically plausible pressure distributions from a single depth image and explores conditional BBDM/LBBDM-based synthesis. While BBDM with ILS yields the strongest metrics, the proposed approach with ILS and WOL achieves higher plausibility with faster training and inference with only a slight decrease in standard scores. Coupled with AttnFnet, it points to a lightweight, non-invasive, real-time patient monitoring system.
@@ 7 Weight Optimation Loss
where $p_{i}$ is the pressure at $i^{th}$ taxel with total $N$ sensors, $A_{i}$ is the area of the $i^{th}$ sensor, and $g$ is gravitational acceleration.
@@
$L_{\text{WOL}}=\|m-\hat{m}\|=\left\|\sum_{i=1}^{N}\frac{p_{i}A_{i}}{g}-\sum_{i=1}^{N}\frac{\hat{p}_{i}\hat{A}_{i}}{g}\right\|$ (20)
@@
$\mathcal{L}_{WOL}=\frac{A_{\text{tot}}}{g}\left\|\sum_{i=1}^{N}(p_{i}-\hat{p}_{i})\right\|$ (21)
@@
$\mathcal{L}_{WOL}\leq\frac{A_{\text{tot}}}{g}\sum_{i=1}^{N}\lVert p_{i}-\hat{p}_{i}\rVert$ (22)
Therefore, We also can reduce BM error though Eq. 23 which is more stable (Fig. 5) but it can’t reduce mass plausibility when low frequency pressure ambiguity in actual distributions:
$\mathcal{L}_{WOL2}=\sum_{i=1}^{N}\lVert p_{i}-\hat{p}_{i}\rVert$ (23)
@@ 8 Effects of Data Normalization and Pre-processing
Figure 6: Comparison of BM estimation accuracy under different data normalization and pre-processing conditions. The scatter plot correlates predicted BM (derived from generated pressure distributions) with actual measurements (red dotted identity line). Blue points represent predictions from the proposed method trained on $8\,\mathrm{k}$ real samples. Notation: g_norm = global normalization; i_norm = individual normalization; $\sigma$ = Gaussian filter standard deviation (e.g., $\sigma=0.5$ denotes filtering with kernel width 0.5); GT denotes BM calculated from ground truth pressure distributions.
@@
Model AttnFnet AttnFnet (small) Openai U-Net Openai U-Net
Pretrained Network - - - AttnFnet (small)
@@
The choice of data normalization strategy significantly impacts the model’s predictive performance for pressure distribution. Normalizing each pressure map individually is problematic, as pressure values are co-dependent on external factors like mattress softness. This approach can obscure the relationship between body mass and pressure features, leading to a decline in performance, as illustrated in Fig. 6. While prior work by Clever et al. [8] addressed this by normalizing pressure with body mass, their method required an seperate network ”Betanet” to predict body mass for denormalization. This network only infers body-mass and body-height and does not build any feature body parameter relationship.
Our approach employs a simpler global pressure normalization, where all pressure distributions are scaled into a [0, 1] range based on a single maximum value from the entire dataset. This method proves to be both plausible and accurate, outperforming baselines without requiring seperate network for body parameters. However, a challenge arises from pre-processing steps like applying a Gaussian filter. While filtering smooths pressure distributions for better approximation (Fig. 6), it can alter the pressure range differently across datasets by reducing low frequency high pressure spike to smoothen the values. For instance, extensive filtering on our real dataset reduces its pressure range to (0, 56)kPa, creating a mismatch with our synthetic data (0, 101)kPa. This discrepancy complicates the generation of valid samples when using WOL.

### v2

# Improving the Plausibility of Pressure Distributions Synthesized from Depth Image through Generative Modeling

@@ Abstract
Monitoring contact pressure in hospital beds is essential for preventing pressure ulcers and enabling real-time patient assessment. Current methods can predict pressure maps but often lack physical plausibility, limiting clinical reliability. This work proposes a framework that enhances plausibility via Informed Latent Space (ILS) and Weight Optimization Loss (WOL) with conditional generative modeling to produce high-fidelity, physically consistent pressure estimates. This study also applies diffusion based conditional Brownian Bridge Diffusion Model (BBDM) and proposes training strategy for its latent counterpart Latent Brownian Bridge Diffusion Model (LBBDM) tailored for pressure synthesis in lying postures. Experiment results shows proposed method improves physical plausibility and performance over baselines: BBDM with ILS delivers highly detailed maps at higher computational cost and large inference time, whereas LBBDM provides faster inference with competitive performance. Overall, the approach supports non-invasive, vision-based, real-time patient monitoring in clinical environments.
@@ 1 Introduction
A major limitation of prior methods is their weak modeling of how pressure depends on human attributes (mass, height, gender). The proposed ILS addresses this by explicitly coupling these factors, producing more plausible and adaptive pressure maps. A practical advantage is the model’s body parameter adaptation: for the same subject’s depth image, adjusting mass, height, or gender yields distinct, physically consistent pressure distributions. This adaptability is clinically useful when patient characteristics change over time, since only the conditioning inputs need updating rather than retraining the model.
This study also introduces pressure synthesis via a conditional Brownian Bridge Diffusion Model (BBDM) and Latent Brownian Bridge Diffusion Model (LBBDM) [23]. We employed a two-stage training strategy for a conditional LBBDM, tailored for pressure synthesis. Additionally, a VQGAN model [42] trained on a distinct objective is integrated with the LBBDM for pressure generation. To see the impact of ILS in the diffusion process, we have incorporated ILS into the denoising model and into the pretrained AttnFnet network when using conditional LBBDM.
We evaluate the proposed method on the publicly available Systematic Lying Postures (SLP) dataset [27] and compare it against established baselines [8, 28].
@@ 2 Related Work
Conventionally, nurses reposition bedridden patients every two hours to reduce skin hydration, minimize prolonged mechanical load, and redistribute pressure [29]. As this repositioning frequency dates back to World War II [11], it remains a standard method for preventing pressure injuries caused by mechanical boundary conditions. However, studies by Peterson et al. [33, 32] suggest that the current repositioning protocol requires improvement, as poor pressure redistribution can still occur despite the scheduled turning.
@@
Building on these prior concepts, the proposed study introduces a novel approach for inferring plausible pressure distributions through conditional generative modeling. This study proposes the inclusion of anthropometric information to build interdependencies via ILS, and further improving pressure plausibility through WOL when predicting distributions in pixel space. Furthermore, in our work, we synthesized pressure distributions using conditional BBDM, evaluated the impact of ILS during diffusion denoising BBDM + ILS, and introduced a two-stage training approach for a conditional LBBDM tailored for pressure synthesis. To the authors’ knowledge, no prior work has applied diffusion-based modeling for predicting pressure distributions.
@@
Although Betanet predicts height and mass, this component remains independent of BPBnet and BPWnet, and does not establish any interrelationships between pressure and anthropometric attributes. Similarly, the approach proposed by Manavar et al. [28] lacks a mechanism to model such dependencies.
This study addresses these challenges by proposing an informed mechanism to incorporate plausibility into pressure inference through ILS and WOL. Additionally, ILS is integrated into the denoising process of the BBDM during training and into the pretrained network during LBBDM training to investigate its influence on pressure plausibility in diffusion denoising.
@@ 3.3 Training Strategy
This study employs conditional generative modeling approaches such as cGAN, BBDM, and LBBDM to synthesize pressure distributions.
We have used the encoder-decoder model architecture introduced by Manavar et al. [28] when predicting pixels. For diffusion model training, we use OpenAI’s improved U-Net as the denoising model, consistent with the setup by Li et al. [23].
@@
Finally, we employed BBDM in the latent space to evaluate computational performance improvements alongside the quality of generated pressure distributions. The AttnFnet model was trained in a self-reconstruction manner using ILS to encode the relationships among mass, height, and gender during image reconstruction. Skip connections in the AttnFnet architecture were removed to minimize the influence of the encoder during the decoding process. To reduce the size of the denoising model, a second bottleneck layer was introduced after the bottleneck layer to decrease the feature channel dimensionality (see supplementary Fig. 7).
@@
This study also utilized an unconditionally trained VQGAN [42] model on the CelebHQ dataset [21], in a manner consistent with the BBDM paper [23] and compared results.
@@ 4 Evaluation
LBBDM proposed pretraining 0.6950 0.8510 0.7338 0.000349 34.93
@@
Manavar et al. [28] AttnFnet-11M 0.6793 0.8847 0.4323 0.000264 36.3225
Ours AttnFnet-11M 0.6451 0.8795 0.6979 0.000270 36.1675
@@
LBBDM proposed pretraining 0.7068 0.8507 0.9406 0.000313 35.4798
$203\,\mathrm{k}$ training samples
Manavar et al. [28] AttnFnet-11M 0.6580 0.8860 0.8014 0.000262 36.4592
Ours $\gamma=0.0182$ AttnFnet-11M 0.6539 0.8942 0.7993 0.000247 36.7770
BBDM Openai U-Net 0.6306 0.8901 0.4267 0.000361 34.9244
BBDM + ILS Openai U-Net 0.4146 0.8630 0.3821 0.000431 34.18
Table 1: Test‑set comparison of MPPA, MSSIM, MFID, MSE, and MPSNR across methods at three training sample sizes. AttnFnet-143M and AttnFnet-11M describes AttnFnet model with 143 million and 11 million parameters respectively.
@@ 4.1 Dataset
The Systematic Lying Postures (SLP) dataset [27] contains 102 healthy participants under three cover conditions and 45 unique poses, grouped into three sleeping postures (supine, left lateral, and right lateral). Cover conditions include uncovered poses, poses covered by a 1 mm blanket, and poses covered by a 3 mm blanket.
@@ 4.2 Data Analysis
Ours $\gamma=0.0182$ 0.6406 0.8814 0.8819 0.8013
BBDM 0.7990 1.077 1.0979 0.9914
BBDM + ILS 1.0031 1.2541 1.2926 1.1833
@@
All models were evaluated under three different training conditions: (1) a sample size of $2745$ ($3\,\mathrm{k}$) using only uncovered postures from real data; (2) a sample size of $8235$ ($8\,\mathrm{k}$) including cover conditions from real data; and (3) a sample size of $203225$ ($203\,\mathrm{k}$) combining both real and synthetic data with all postures and covered conditions. All results are reported at epoch 100, unless stated otherwise.
Standard evaluation metrics for 2D images were used to assess pressure generation quality: Mean Pixel Prediction Accuracy (MPPA) [28], Mean Structural Similarity Index (MSSIM) [40], Mean Fréchet Inception Distance (MFID) [37], MSE, and Mean Peak-Peak Signal-to-Noise Ratio (MPSNR) [13]. Fréchet inception distance (FID) was calculated with the original Inception-V3 model weights [16]. Table 1 presents metric scores for the proposed methods and baselines on the test dataset. In Tab. 1, BBDM with ILS outperformed all methods when trained on $8\,\mathrm{k}$ samples, while our method with $\gamma=0.0182$ outperformed in $203\,\mathrm{k}$ training samples.
When images were rescaled from the normalized space, MSE was recalculated in pressure space as shown in Tab. 2. Consistent trends were observed: the proposed method displayed increased MSE in pressure space, but error rate declined when setting $\gamma=0.0182$. Note that pressure ranges differ between real and synthetic data, and setting $\gamma=1.0$ prevents the model from generating valid samples.
@@
Participant BMs were calculated from actual and predicted pressure profiles using Eq. 9. Table 3 reports mean absolute BM differences among measured BMs ($M_{\text{meas}}$), BMs calculated from the ground truth pressures ($M_{\text{GT}}$), and from predicted pressures ($M_{\text{pred}}$). Further, BM plots in Fig. 3 enables detailed comparisons.
@@ 5 Ablation Study
Table 5 demonstrates that combining ILS and WOL yields more plausible pressure distributions with a slight increase in MSE. Table 6 presents the effect of the WOL weight $\gamma$ on MSE and BM Mean Absolute Error (MAE): reducing $\gamma$ initially reduces the MSE but, increases the BM error.
@@
Table 5: Effect of the ILS and WOL components on prediction error: MSE of pressure maps and BM MAE. Model was trained using $8\,\mathrm{k}$ real samples.
$\gamma$ values MSE ($kPa^{2}$) BM MAE
@@
Table 6: Effect of the WOL weight $\gamma$ on pressure MSE and BM MAE. Model was trained using $8\,\mathrm{k}$ real samples.
Sampling steps MSE ($kPa^{2}$) BM MAE FID
@@
Table 7: Effect of diffusion sampling steps on pressure MSE, BM MAE, and FID; model was trained using $8\,\mathrm{k}$ real samples.
@@ 6 Discussion and Future Work
According to Tab. 1 and Tab. 2, BBDM+ILS and BBDM perform best when trained on real data only. However, performance degrades when trained on the larger $203\,\mathrm{k}$ dataset. We speculate that these models require more epochs to converge on the larger dataset compared to our method, though we could not validate this during training due to high inference times and limited compute resources. Nevertheless, Fig. 4 shows that BBDM+ILS produces visually superior results compared to the standard BBDM.
A key requirement for using ILS in the model is establishing a prior global context for the features, as ILS cannot generate valid samples during training without it.
@@ Conclusion
This work predicts physically plausible pressure distributions from a single depth image and explores conditional BBDM/LBBDM-based pressure synthesis. While BBDM with ILS yields the strongest metrics, the proposed approach with ILS and WOL achieves higher plausibility with faster training and inference with only a slight decrease in standard scores. Coupled with AttnFnet, it points to a lightweight, non-invasive, real-time patient monitoring system.
@@ 7 Weight Optimation Loss
where $p_{i}$ is the pressure at $i^{th}$ taxel with total $N$ sensors, $A_{i}$ is the area of the $i^{th}$ taxel, and $g$ is gravitational acceleration.
@@
$L_{\text{WOL}}=|m-\hat{m}|=\left|\sum_{i=1}^{N}\frac{p_{i}A_{i}}{g}-\sum_{i=1}^{N}\frac{\hat{p}_{i}\hat{A}_{i}}{g}\right|$ (20)
@@
$\mathcal{L}_{WOL}=\frac{A_{\text{tot}}}{g}\left|\sum_{i=1}^{N}(p_{i}-\hat{p}_{i})\right|$ (21)
@@
$\mathcal{L}_{WOL}\leq\frac{A_{\text{tot}}}{g}\sum_{i=1}^{N}|p_{i}-\hat{p}_{i}|$ (22)
Therefore, reducing BM error is same as minimizing $L_{1}$ norm Eq. 23. It is more stable (Fig. 5) but it can’t reduce mass plausibility when low frequency pressure ambiguity in actual distributions:
$\mathcal{L}_{WOL2}=\sum_{i=1}^{N}|p_{i}-\hat{p}_{i}|=\|\mathbf{p}-\hat{\mathbf{p}}\|_{1}$ (23)
@@ 8 Effects of Data Normalization and Pre-processing
Figure 6: Comparison of BM estimation accuracy under different data normalization and pre-processing conditions. The scatter plot correlates predicted BM (derived from generated pressure distributions) with actual measurements (red dotted line). Blue points represent predictions from the proposed method trained on $8\,\mathrm{k}$ real samples. Notation: g_norm = global normalization; i_norm = individual normalization; $\sigma$ = Gaussian filter standard deviation (e.g., $\sigma=0.5$ denotes filtering with kernel width 0.5); GT denotes BM calculated from ground truth pressure distributions.
@@
Model AttnFnet-143M AttnFnet-11M Openai U-Net Openai U-Net
Pretrained Network - - - AttnFnet-11M
@@
The choice of data normalization strategy significantly impacts the model’s predictive performance for pressure distribution. Normalizing each pressure map individually is problematic, as pressure values are co-dependent on external factors like mattress softness. This approach can obscure the relationship between body mass and pressure features, leading to a decline in performance, as illustrated in Fig. 6. While prior work by Clever et al. [8] addressed this by normalizing pressure with body mass, their method required an seperate network ”Betanet” to predict body mass for denormalization. Betanet only infers BM and body-height and does not build any feature body parameter relationship.
Our approach employs a simpler global pressure normalization, where all pressure distributions are scaled into a [0, 1] range based on a single maximum value from the entire dataset. This method proves to be both plausible and suitable for our approach (Fig. 6). However, a challenge arises from pre-processing steps like applying a Gaussian filter. While filtering smooths pressure distributions for better approximation, it can alter the pressure range differently across datasets by reducing low frequency high pressure spike to smoothen the values. For instance, extensive filtering on our real dataset reduces its pressure range to (0, 56)kPa, creating a mismatch with our synthetic data (0, 101)kPa. This discrepancy complicates the generation of valid samples when using WOL.